NOAA Satellite Information System

Special Bulletins Archive

 

Red dot NOAA is experiencing degradation of NOAA-18 AVHRR products. The degraded products are noticeable since sensing time 21:18 on 30.09.2008. We are investigating our system to try to identify the cause. More information will follow when available.


Red dot The Metop-A HRPT instrument will be will be back to routine operation on Monday,the 29th of September. The instrument will not image during entire orbits. A geographic zone covering the main parts of North of Europe and North of Atlantic has been chosen to ensure a minimum risk of failure. Therefore, for all descending orbits in daylight over Europe, the modulator of the HRPT will be switched on.

The schedule to resume HRPT routine operations will be the following:
- Monday the 29th, pass ~08:00z, SSPA will be switch ON manually, on B-side of Metop-A, using the nominal frequency.
- If this switch ON is successful, the modulator will be switched ON/OFF via Time-Tag commands. Please find hereafter the exact switch ON/OFF time for next week. That will be the start of routine operation of HRPT for Metop-A.

The planning for next week is the following :
Orbit - Switch ON
10097 - 2008/09/29 09:57:52.955 2008/09/29 10:12:53.776
10098 - 008/09/29 11:39:14.475 2008/09/29 11:52:37.556
10099 - 2008/09/29 13:21:56.180 2008/09/29 13:31:41.123
10111 - 2008/09/30 09:36:55.030 2008/09/30 09:51:58.951
10112 - 2008/09/30 11:18:16.551 2008/09/30 11:31:28.165
10113 - 2008/09/30 13:00:41.719 2008/09/30 13:09:24.443
10124 - 2008/10/01 07:36:31.036 2008/10/01 07:47:28.327
10125 - 2008/10/01 09:15:57.052 2008/10/01 09:29:46.708
10126 - 2008/10/01 10:57:18.575 2008/10/01 11:10:18.718
10127 - 2008/10/01 12:39:27.100 2008/10/01 12:48:08.753
10138 - 2008/10/02 07:16:02.138 2008/10/02 07:26:05.882
10139 - 2008/10/02 08:54:59.072 2008/10/02 09:06:51.242
10141 - 2008/10/02 12:18:12.368 2008/10/02 12:28:22.372
10142 - 2008/10/02 14:00:53.863 2008/10/02 14:07:11.758

Orbit - Switch OFF
10110 - 2008/09/30 07:56:55.998 2008/09/30 08:08:32.092
10140 - 2008/10/02 10:36:20.605 2008/10/02 10:49:43.842


Red dot Topic: RESCHEDULED GOES-12 (GOES-East) East/West Maneuver August 20/21, 2008
Message Issued: August 18, 2008, 1830 UTC Satellites Impacted: GOES-12 (GOES-East) and GOES-11 (GOES-West)
Products Impacted: GOES-12 (GOES-East) Imager and Sounder Data and GOES-11 (GOES-West) Imager Data
Date/Time of Initial Impact: August 20, 2008 1936 UTC
Details: A GOES-12 (GOES-East) East-West Station Keeping maneuver will be performed on Thursday, August 21, 2008 at 0159 UTC. This maneuver time was postponed from the original scheduled date of August 21, due to the satellite coverage needed for Tropical Storm Fay. This is the final date available for this maneuver, therefore there is no backup date. On August 20, the following exclusive schedule will be employed to support maneuver operations: (*PLEASE NOTE the scheduled GOES-11 (GOES-West) Full Disk scan period to support the primary GOES-12 (GOES-East) outage during the maneuver.*)

GOES-12 (GOES-East) Maneuver Schedule starts at 1936 UTC on August 20, 2008.
(Southern Hemisphere Scan at 1939 UTC is canceled) From 20/1945 UTC through 21/0114 UTC --
GOES-12 (GOES-East) Routine Imaging and Soundings. From 21/0115 UTC through 21/0144 UTC --
GOES-12 (GOES-East) Full Disk Imaging and Routine Soundings *From 21/0145 UTC through 21/0314 UTC --
No GOES-12 (GOES-East) Imaging or Soundings* **From 21/0130 UTC through 21/0329 UTC --
GOES-11 (GOES-West) Full Disk Imaging** From 21/0315 UTC through 21/1014 UTC --
GOES-12 (GOES-East) Post Maneuver Recovery Including:
* Full Disk Images (19 minute scan) at XX15 and XX45
* CONUS Soundings at XX46
* INR recovery for GOES-12 will run through the GOES-12 KOZ period (see http://www.ssd.noaa.gov/PS/SATS/eclipse.html) 1015 UTC August 21 --
GOES-12 (GOES-East) Routine Scan Operations Resume Expect GOES-12 (GOES-East) imager and sounder gridding offsets for up to 6 hours following the maneuver.


Red dot There will be a planned outage of LRIT on both GOES-11 and GOES-12 today, July 10 at approximately 18:30 UTC.
The duration of the outage will be approximate 10-20 minutes.
The reason for the outage is NOAA will be replacing communications equipment in the system.
New information will be posted when it becomes available.


Red dot Topic: REMINDER GOES-11 (GOES-West) North/South Maneuver June 17, 2008
Message Issued: June 16, 2008, 1350 UTC
Satellites Impacted: GOES-11 (GOES-West) and GOES-12 (GOES-East)
Products Impacted: GOES-11 (GOES-West) Imager and Sounder Data and GOES-12 (GOES-East) Imager Data
Date/Time of Initial Impact: June 17, 2008 0147 UTC
Details: A GOES-11 (GOES-West) North-South Station Keeping maneuver will be performed on Tuesday, June 17, 2008 at 0858 UTC. The backup date for this maneuver is Thursday, June 19, 2008. If the NWS regions have any operational concerns, please contact the NCEP SDM (301-763-8298) at least 6.5 hours prior to the maneuver time.

On June 17, the following exclusive schedule will be employed to support maneuver operations:
(*PLEASE NOTE the scheduled GOES-12 (GOES-East) Full Disk scan period to support the primary GOES-11 (GOES-West) outage during the maneuver.*) GOES-11 (GOES-West) Maneuver Schedule starts at 0147 UTC on June 17, 2008. (Southern Hemisphere scan at 0151 UTC is canceled) From 0200 UTC through 0729 UTC -- GOES-11 (GOES-West) Routine Imaging and Soundings.
From 0730 UTC through 0759 UTC -- GOES-11 (GOES-West) Full Disk imaging and routine soundings.
*From 0800 UTC through 1029 UTC -- No GOES-11 (GOES-West) Imaging or Soundings*
**From 0815 UTC through 1044 UTC -- GOES-12 Full Disk Imaging** From 1030 UTC through 1859 UTC -- GOES-11 (GOES-West) Post Maneuver Recovery Including:
* Full Disk Images (19 minute scan) at XX00 and XX30
* CONUS Soundings at XX01
1900 UTC June 17 -- GOES-11 (GOES-West) Routine Scan Operations Resume Expect GOES-11 (GOES-West) imager and sounder gridding offsets for up to 6 hours following the maneuver.


Red dot Topic: (GOES-South America) Navigation
Problem Message Issued: June 11, 2008, 1855 UTC
Satellites Impacted: GOES-10 Products Impacted:
Imager products from GOES-10
Date/Time of Initial Impact: June 10, 2008 2005 UTC


Red dot Topic: GOES-11 (GOES-West) North/South Maneuver June 17, 2008
Message Issued: June 9, 2008, 1145 UTC
Satellites Impacted: GOES-11 (GOES-West) and GOES-12 (GOES-East)
Products Impacted: GOES-11 (GOES-West) Imager and Sounder Data and GOES-12 (GOES-East) Imager Data
Date/Time of Initial Impact: June 17, 2008 0147 UTC
Details: A GOES-11 (GOES-West) North-South Station Keeping maneuver will be performed on Tuesday, June 17, 2008 at 0858 UTC. The backup date for this maneuver is Thursday, June 19, 2008. If the NWS regions have any operational concerns, please contact the NCEP SDM (301-763-8298) at least 6.5 hours prior to the maneuver time.

On June 17, the following exclusive schedule will be employed to support maneuver operations:
(*PLEASE NOTE the scheduled GOES-12 (GOES-East) Full Disk scan period to support the primary GOES-11 (GOES-West) outage during the maneuver.*)
GOES-11 (GOES-West) Maneuver Schedule starts at 0147 UTC on June 17, 2008.
(Southern Hemisphere scan at 0151 UTC is canceled)
From 0200 UTC through 0729 UTC -- GOES-11 (GOES-West) Routine Imaging and Soundings.
From 0730 UTC through 0759 UTC -- GOES-11 (GOES-West) Full Disk imaging and routine soundings.
*From 0800 UTC through 1029 UTC -- No GOES-11 (GOES-West) Imaging or Soundings*
**From 0815 UTC through 1044 UTC -- GOES-12 Full Disk Imaging**
From 1030 UTC through 1859 UTC -- GOES-11 (GOES-West) Post Maneuver Recovery Including:
* Full Disk Images (19 minute scan) at XX00 and XX30
* CONUS Soundings at XX01

1900 UTC June 17 -- GOES-11 (GOES-West) Routine Scan Operations Resume Expect GOES-11 (GOES-West) imager and sounder gridding offsets for up to 6 hours following the maneuver.


Red dot The ESPC SATEPSPROD systems are now receiving and processing data to produce products, with the exception of SATEPSPROD1 which processes the following products; ASOS Imager,GCIP,GSIP DPI ABBA Products, ABBA Transfers,Volcano, and HYDRO-Estimator.

Work on SATEPSPROD1 is nearly completed and ESPC expects these products will be available very soon.


Red dot Details of the Outage: There will be a scheduled outage of many ESPC systems today, Thursday June 5, 2008 1430-1600Z. This outage is necessary to replace a failed controller on the ESPC SAN. Most ESPC products will be affected by this outage. However, direct readout users will not be impacted, nor will imagery delivered to AWIPS. June 5 was declared a critical weather day, this outage is still scheduled.
Data Affected by the Outage: Most ESPC Products.
Date and Time of the Outage: 06/05/2008, 14:30 UTC, 1130 AM EST
Length of the Outage: Approx. 1 hour and 30 minutes


Red dot Topic: GOES-10 (GOES South America) East/West Maneuver June 10, 2008
Message Issued: June 4, 2008, 1415 UTC
Satellites Impacted: GOES-10
Products Impacted: GOES-10 Imager and Sounder Data
Date/Time of Initial Impact: June 10, 2008 1233 UTC

A GOES-10 East-West Station Keeping maneuver will be performed on Tuesday, June 10, 2008 at 1813 UTC. There is no back-up day scheduled for this maneuver.
On June 10, the following exclusive schedule will be employed to support maneuver operations:
GOES-10 Maneuver Schedule starts at 1233 UTC on June 10, 2008.
(GOES-10 South America scan at 1245 UTC is canceled.)
From 1257 UTC through 1644 UTC -- GOES-10 Routine Imaging and Soundings.
From 1645 UTC through 1714 UTC -- GOES-10 Full Disk Imaging and Soundings.
From 1715 UTC through 1754 UTC -- GOES-10 South America Imaging and Soundings.
*From 1757 UTC through 1914 UTC -- No GOES-10 Imaging or Soundings*
From 10/1915 UTC through 11/0205 UTC -- GOES-10 Post Maneuver Recovery
Including:
* Americas Scans (19:30 minute scan) at XX15 and XX45
* Abbreviated South America 1-4 Soundings occur at XX31

0215 UTC June 11 -- GOES-10 Routine Scan Operations Resume
Expect GOES-10 imager and sounder gridding offsets for up to 6 hours following the maneuver.


Red dot Topic: REMINDER GOES-12 (GOES-East) East/West Maneuver June 3, 2008
Message Issued: June 2, 2008, 1745 UTC
Satellites Impacted: GOES-12 (GOES-East) and GOES-11 (GOES-West)
Products Impacted: GOES-12 (GOES-East) Imager and Sounder Data and GOES-11 (GOES-West) Imager Data
Date/Time of Initial Impact: June 3, 2008 1405 UTC
Details:
A GOES-12 (GOES-East) East-West Station Keeping maneuver will be performed on Tuesday, June 3, 2008 at 2021 UTC. For National Weather Service operations: The backup date for this maneuver is Thursday, June 5, 2008. If the NWS regions have any operational concerns, please contact the NCEP SDM (301-763-8298) at least 6.5 hours prior to the maneuver time.

On June 3, the following exclusive schedule will be employed to support maneuver operations: (*PLEASE NOTE the scheduled GOES-11 (GOES-West) Full Disk scan period to support the primary GOES-12 (GOES-East) outage during the maneuver.*)

GOES-12 (GOES-East) Maneuver Schedule starts at 1405 UTC on June 3, 2008. (Southern Hemisphere Scan at 1409 UTC is canceled)

From 1415 UTC through 1944 UTC -- GOES-12 (GOES-East) Routine Imaging and Soundings.
From 1045 UTC through 2014 UTC -- GOES-12 (GOES-East) Full Disk Imaging and Routine Soundings
*From 2015 UTC through 2114 UTC -- No GOES-12 (GOES-East) Imaging or Soundings*
**From 2000 UTC through 2129 UTC -- GOES-11 (GOES-West) Full Disk Imaging**
From 3/2115 UTC through 4/0414 UTC -- GOES-12 (GOES-East) Post Maneuver Recovery Including:
* Full Disk Images (19 minute scan) at XX15 and XX45
* CONUS Soundings at XX46
0415 UTC June 4 -- GOES-12 (GOES-East) Routine Scan Operations Resume
Expect GOES-12 (GOES-East) imager and sounder gridding offsets for up to 6 hours following the maneuver.


Red dot Details of the Outage: ESPC is not receiving the HRPT data. System programmer is working on the problem.
Data Affected by the Outage: HRPT (High Resolution Picture Transmission) data
Date and Time of the Outage: 05/27/2008, 1700 UTC, 01:00PM EDT
Length of the Outage: Until further norice


Red dot Topic: REMINDER GOES-11 (GOES-West) East/West Maneuver April 29, 2008
Message Issued: April 29, 2008, 0200 UTC
Satellites Impacted: GOES-11 (GOES-West) and GOES-12 (GOES-East)
Products Impacted: GOES-11 (GOES-West) Imager and Sounder Data and GOES-12 (GOES-East) Imager Data
Date/Time of Initial Impact: April 29, 2008 0547 UTC

Details:
A GOES-11 (GOES-West) East-West Station Keeping maneuver will be performed on Tuesday, April 29, 2008 at 1103 UTC. The backup date for this maneuver is Thursday, May 1, 2008.
On April 29, the following exclusive schedule will be employed to support maneuver operations:
(*PLEASE NOTE the scheduled GOES-12 (GOES-East) Full Disk scan period to support the primary GOES-11 (GOES-West) outage during the maneuver.*)
GOES-11 (GOES-West) Maneuver Schedule starts at 0547 UTC on April 29, 2008.
(Southern Hemisphere scan at 0551 UTC is canceled)
From 0600 UTC through 0929 UTC -- GOES-11 (GOES-West) Routine Imaging and Soundings.
From 0930 UTC through 1051 UTC -- GOES-11 (GOES-West) Full Disk imaging (30 minute) and routine soundings.
From 1052 UTC through 1159 UTC -- No GOES-11 (GOES-West) Imaging or Soundings
From 1045 UTC through 1214 UTC -- GOES-12 Full Disk (30 minute) Imaging
From 1200 UTC through 1859 UTC -- GOES-11 (GOES-West) Post Maneuver Recovery Including:
* Full Disk Images (19 minute scan) at XX00 and XX30
* CONUS Soundings at XX01
1900 UTC April 29 -- GOES-11 (GOES-West) Routine Scan Operations Resume
Expect GOES-11 (GOES-West) imager and sounder gridding offsets for up to 6 hours following the maneuver.


Red dot Details of the Outage: In-plane manoeuvre 2008-04-24. More information to follow. GOME - to manoeuvre timeline. Sensing Stop 2008-04-24 11:30 - 16:30 SEM - to standby 2008-04-24
Data Affected by the Outage: Polar data
Date and Time of the Outage: 4/24/2008, Unknown at this time
Length of the Outage: Until further norice


Red dot UPDATE - GOES-12 (GOES-EAST) Will be restored to normal operation at 0034 UTC on J/D- 109
Details of the Outage: GOES-12 (GOES-EAST) will be placed into maneuver recovery schedule.
Full Disk Images (19 minute scan) at XX15 and XX45
Data Affected by the Outage: GOES-12
Date and Time of the Outage: 1615 UTC
Length of the Outage: Up to 24 Hours


Red dot Details of the Outage: GOES-12 (GOES-EAST) will be placed into maneuver recovery schedule.
Full Disk Images (19 minute scan) at XX15 and XX45
Data Affected by the Outage: GOES-12
Date and Time of the Outage: 1615 UTC
Length of the Outage: Up to 24 Hours


Red dot Data Outage: The GOES-12 Grids are off. OSO (SOCC) is investigating the problem.
Data Affected by the Outage: Imager Data
Date and Time of the Outage: 04/17/2008 0240 UTC
Length of the Outage: *Until further Notice


Red dot ESPC is now receiving Metop-2 (Meteorological Operational Satellite) data. The problem was SKYX needed to be reloaded. The problem has been resolved. ESPC hasn't received the following FRAC datasets.
NCS.FRAC.M2.B774041.SV EST. due in 15:05 UTC
NCS.FRAC.M2.B774243.SV EST. due in 18:28 UTC
NCS.FRAC.M2.B774344.SV EST. due in 20:09 UTC
NCS.FRAC.M2.B774445.SV EST. due in 21:50 UTC
Data Affected by the Outage: AVHRR, HIRS, etc.
(Same type of data that is generated from the POLAR data.)
Date and Time of the Outage: 4/16/2008, 1505 UTC
Length of the Outage: 4/16/2008, 1505 UTC to 2205 UTC, 7 Hours


Red dot ESPC has not received the following METOP-2/A orbits:
NCS.FRAC.M2. B774041.SV at 15:05 UTC and orbit NCS.FRAC.M2.B774243.SV at 1828 UTC.
The problem is being investigated.
Data Affected by the Outage: METOP-2/A Data Date and Time of the Outage: 04/16/2008 1505 UTC Length of the Outage: Until further notice.


Red dot Topic: CORRECTION GOES-12 (GOES-East) East/West Maneuver April 15, 2008
Message Issued: April 15, 2008, 1115 UTC
Satellites Impacted: GOES-12 (GOES-East) and GOES-11 (GOES-West)
Products Impacted: GOES-12 (GOES-East) Imager and Sounder Data and GOES-11 (GOES-West) Imager Data
Date/Time of Initial Impact: April 15, 2008 1305 UTC
Details: Correction issued for the length of the GOES-12 (GOES-East) outage and the length of the GOES-11 (GOES-West) Full Disk schedule
A GOES-12 (GOES-East) East-West Station Keeping maneuver will be performed on Tuesday, April 15, 2008 at 1916 UTC. For National Weather Service operations: The backup date for this maneuver is Thursday, April 17, 2008. If the NWS regions have any operational concerns, please contact the NCEP SDM (301-763-8298) at least 6.5 hours prior to the maneuver time.

On April 15, the following exclusive schedule will be employed to support maneuver operations: (*PLEASE NOTE the scheduled GOES-11 (GOES-West) Full Disk scan period to support the primary GOES-12 (GOES-East) outage during the maneuver.*)

GOES-12 (GOES-East) Maneuver Schedule starts at 1305 UTC on April 15, 2008.
(Southern Hemisphere Scan at 1309 UTC is canceled)
From 1315 UTC through 1829 UTC -- GOES-12 (GOES-East) Routine Imaging and Soundings.
*From 1830 UTC through 2014 UTC -- No GOES-12 (GOES-East) Imaging or Soundings*
**From 1830 UTC through 2029 UTC -- GOES-11 (GOES-West) Full Disk Imaging**
From 15/2015 UTC through 16/0314 UTC -- GOES-12 (GOES-East)
Post Maneuver Recovery Including:
* Full Disk Images (19 minute scan) at XX15 and XX45
* CONUS Soundings at XX46
0315 UTC April 16 -- GOES-12 (GOES-East) Routine Scan Operations Resume Expect GOES-12 (GOES-East) imager and sounder gridding offsets for up to 6 hours following the maneuver.


Red dot Topic: REMINDER GOES-12 (GOES-East) East/West Maneuver April 15, 2008
Message Issued: April 14, 2008, 1340 UTC
Satellites Impacted: GOES-12 (GOES-East) and GOES-11 (GOES-West)
Products Impacted: GOES-12 (GOES-East) Imager and Sounder Data and GOES-11 (GOES-West) Imager Data
Date/Time of Initial Impact: April 15, 2008 1305 UTC
Details: A GOES-12 (GOES-East) East-West Station Keeping maneuver will be performed on Tuesday, April 15, 2008 at 1916 UTC.


Red dot OSO/SOCC has recovered GOES-12. At around 1730 UTC (1:30 pm EDT). OSO will begin imaging and sounding on GOES-12. The imager will be placed into a Rapid Scan schedule in support of severe weather monitoring in the Southern U.S. However, the RSO might result in the degradation of the Image and Navigation Registration (INR). There may be a gridding offset for several hours. If the image navigation becomes a problem to users, NESDIS and NWS will coordinate on setting the GOES-12 image back to a normal schedule in order to take additional landmarks to correct the INR. GOES-11 imager will be placed back into routine schedule.


Red dot NOAA are currently experiencing an anomaly on GOES-12 (April 9, 2008). We had a large attitude disturbance in the pitch axis. GOES-11 is currently providing full disk imaging coverage and will remain in that mode until full operations are restored on GOES-12. At this point there is no estimate on when GOES-12 will return to service. At this time, the services: EMWIN, LRIT and DCS are still on.


Red dot Topic: Official notice declaring channel 4 unusable
Date/Time(UTC) Message Issued: 26 March 2008, 1630 UTC
Satellite(s) Involved: NOAA-16
Instrument(s) Involved: AMSU-A1
Product(s) Involved/Affected: Channel 4 Products
Date/Time(UTC) of Initial Implementation: 4 March 2008 Time N/A
Details/Specifics of Change: The NOAA-16 AMSU-A1 Channel.4 has been showing indications of it malfunctioning since January 2007. However, between 25 February and 4 March, the problem increased to the point that over half the data became unusable. Its NEdT increased to >5K (nominal value ~0.2 K). The channel is deemed unusable and is not expected to recover. Other channels are still functional with some limitations.


Red dot Metop-A AVHRR dissemination resumed with the instrument sensing time of 15:14 at 16:52.
Data Affected by the Outage: AVHRR (Advanced Very High Resolution Radiometer)
Date and Time of the Outage: 3/19/2008, 0100 UTC, 9:00 PM, EDT
Length of the Outage: From 3/19/2008, 0100 UTC, 9:00 PM, EDT to 3/21/2008, 16:52 UTC, 12:52 PM, EDT 39 hours and 52 minutes


Red dot Details of the Outage:
The recovery efforts for PLSOL have begun, a little earlier than scheduled. At, 3:38p.m. EDT, the NOAA interface unit was turned on. This provided the ability to view temperature data from the instruments. Engineers at SOCC were able to view the temperatures of the NOAA instruments and feel that they are nominal and activation can proceed.

We are still looking at Friday, 3/21/08, to complete the activation of the prime NOAA instruments. EUMETSAT instruments activation should be completed on Saturday. This may change based on the performance and responses to the activation sequences.

METOP-A data will not be available until further notice.

Data Affected by the Outage:
A-DCS/2, AMSU-A, ASCAT, AVHRR, GRAS, GOME HIRS, IASI, MHS, S&R, SEM

Date and Time of the Outage:
03/19/2008, 0100 UTC 09:00 P.M.


Red dot METOP-A data will not be available until further notice. Data Affected by the Outage:  A-DCS/2, AMSU-A, ASCAT, AVHRR, GRAS, GOME HIRS, IASI, MHS, S&R, SEM. Date and Time of the Outage: 03/20/2008, 0100 UTC 09:00 P.M. Length of the Outage: Unknown


Red dot On Sunday, 2 Mar 08 at 20:35z there was a drop in the NOAA-17 STX-1 (1698Mhz) output power from 5.1W to 2.29W. STX-1 is still transmitting HRPT successfully. NOAA has seen two (2) sudden decreases in power within 7 days. On 27 Feb, this transmitter decreased in power output from 5.6W to 5.1W. HRPT transmissions remain successful at this time. Transmitters remaining on NOAA-17 are STX-2 (Omni/Directional -1702.5MHz) and STX-4 (2247.5MHz) both currently used for PBKS.


Red dot Due to the failure of STX-3 on NOAA-17 SOCC has had to make configuration changes to transmitters used for HRPT and playbacks. NOAA-18 and NOAA-17 are affected by these changes. These changes go into effect on 26 Feb 08 (JD057) on the revs (orbits) and UTC times noted in the table. This configuration will remain indefinitely.

Constellation transmitter changes are provided below.

*Spacecraft*
*HRPT Broadcast*
*Playback transmitters*
*Time of change*

NOAA-17*
STX-1
1698.0 Mhz
STX-2
STX-4
Rev:29485
Time:01:01:15z

NOAA-18*
STX-3
1707.0 Mhz
STX-1
STX-2
Rev:14265
Time:00:35:30z

NOAA-15
STX-2 (OMNI) 1702.5 Mhz
STX-4
No change

NOAA-16
STX-2
1702.5 Mhz/LCP
STX-1
No change

Please note that NOAA-17 has been using STX-2 since the failure of its STX-3 on 14 Feb 08. This allowed NOAA/SOCC to provide as much data as possible to users while schedules were being changed to this final configuration.


Red dot The SOCC is going to do another test for the 2.66Mbps data rate for POES transfers. At the completion of scheduled post pass data, the following passes will be configured for the 2.66Mbps rate, and the GAC and LAC data will be transferred to ESPC for data comparison.

Thursday, February, 21, 2008: N-17 29422/W 080221 1526z 1541z

SOCC will alert WCDAS Operations during this pass to prepare for the configuration change to 2.66 immediately following the normal post pass transfers.

Friday, February 22, 2008: N-17 29436/W 080222 1503z 1518z


Red dot At 4pm local time satellite operations will switch from transmitter #4 2247.5 MHz to transmitter #2 1702.5 MHz, but will use the emergency antenna. The emergency antenna is an omni RHC polarization. This should allow all HRPT site to access the data. There is some concern that there may be an issue with collecting global data at the CDA sites on transmitter #1 1698MHz due to the close proximity of transmitter #2 frequency. Tests will be conducted this afternoon. Ultimately the plan is to move NOAA-17 HRPT to transmitter #1 and NOAA-18 HRPT to transmitter #3 by February 26th.


Red dot A GOES-11 (GOES-West) East-West Station Keeping maneuver will be performed on Tuesday, February 19, 2008 at 1130 UTC. The backup date for this maneuver is Thursday, February 21, 2008.


Red dot Today, Thursday, 14 Feb 2008, at 14:49z (9:49 a.m.east coast local time) the HRPT stream has been moved to transmitter #4 2247.5MHz. We will continue to broadcast HRPT on that frequency until further notice. It is important to note that in the near future users can expect to see a change in HRPT broadcast frequencies for NOAA-18 and NOAA-17. Engineers continue to evaluate all data types and transmitter options across the constellation. SOCC engineers have confirmed failure of transmitter #3 on NOAA-17 (1707MHz) that was used to broadcast HRPT data.

HRPT users are requested to provide feedback on the impact of moving the broadcast to transmitter 4 to their HRPT reception Please send emails to paul.seymour@noaa.gov . Thank you.


Red dot This a notice on an HRPT outage on NOAA 17. On February 13th or JD045 at 00:45z STX3 stopped transmitting HRPT.Investigations are underway. Estimated time to recovery is unknown at this time.


Red dot NOAA/NESDIS/OSDPD/OSO is moving to upgrade the domsat links from the CDAS's that carry the POES GAC, LAC and HRPT data.
The data rate will be 2.66MBPS which is twice the current rate.
There will be some tests of some test passes next Monday, the 11th.
They will most likely be in the 14z-16z time frame. Here are the exact revs and times..
AOS   FRT   FCT   LRT   LCT   AOS
17 29279 W 22.89 080211 141830 141910 142000 143050 143130 143220
17 29280 W 39.81 080211 155800 155900 155920 161120 161200 161240


Red dot We have received the follow notice from EUMETSAT: Operational Services derived from Metop-A are suspended until further notice. Additional information will be provided when it is received from EUMETSAT.


Red dot During rev 28899 on January 15, the cross strapping unit (XSU) on NOAA-17 will be switched to the backup side. This should occur at approximately 21:18z. There will be an outage of HRPT during the contact while the transmitters are off. TIP will be configured to beacon BTX#2, 137.77MHz. Users may also see a shift in the instrument analog values as the XSU switch causes clocks to re-syncronize. This is similar to what was observed on NOAA-16. HRPT should be commanded on before loss of signal over the CDA. There will be a gap in the data of approximately one major frame.


Red dot Due to equipment maintenance at Wallops CDA there will be an LRIT service outage today beginning at 10:00am EST. The schedule is as follows:
GOES West outage 15:00z - 18:00z
GOES East outage 18:00z - 21:00z


Red dot GOES-East product processing was successfully switched back to GOES-12 at 17:45 UTC on December 17, 2007. Users of ESPC GOES-East and NOAAPORT/AWIPS users should be seeing 30 minute GOES-East imagery at XX:15 and XX:45. The GOES Sounder products should show up soon on NOAAPORT. GOES-12 will resume the normal schedule at 20:45 UTC. See schedules at: http://www.ssd.noaa.gov/PS/SATS/GOES/EAST/sched.html (web page will be updated shortly)

**GOES-EAST users receiving GOES-10 directly from the satellite should re-point their antennas to GOES-12. We plan to release GOES-10 back into GOES-South America schedule at or before 2100 UTC** Users of auxiliary services such as LRIT, DCS, or EMWIN should retrieve the signal from GOES-12 (we will drop the retransmission through GOES-10 today)


Red dot GOES-10 will continue to run the GOES-East schedule through next Monday. All auxiliary data, EMWIN, DCS and LRIT, will remain on GOES-12. It is anticipated the GOES-12 will return to normal operations on or about December 17th.


Red dot At 11:47z yesterday morning, engineers performed a E/W maneuver on GOES-12 to start a western drift of 0.5 degrees per day. Following the maneuver, INR recovery was started. For INR recovery, the GVAR broadcast on GOES-12 was switched from GOES-10 data to GOES-12 data. Next Monday, there will be one more E/W maneuver on GOES-12 to stop the western drift. GOES-10 will continue to run the GOES-East schedule through next Monday. All auxiliary data, EMWIN, DCS and LRIT, will remain on GOES-12.


Red dotThe following are updates on GOES-10 and GOES-12:
1. GOES-10: GOES-10 will continue to operate in GOES-EAST mode until Monday, December 17.
2. GVAR will be broadcast from GOES-10 and through GOES-12 (**through GOES-12 until Tuesday December 11 -- see GOES-12 update below**).
3. Ancillary services such as DCS, LRIT, and EMWIN will be broadcast from GOES-10 and GOES-12 until next week.
4. GOES-12: GOES-12 was successfully returned to normal on orbit mode as of Friday evening, December 7, 2007.
5. On Tuesday, December 11, NESDIS will perform a maneuver to move GOES-12 back to its nominal position of 75 degrees West.
6. The maneuver will take place starting at 1146 UTC on Tuesday, December 11.
7. Following the maneuver on Tuesday, December 11, users receiving GVAR from GOES-12 will see the GVAR rebroadcast from GOES-10 dropped in favor of the GOES-12 imager/sounder GVAR startup.


Red dot GOES-12 remains in a power positive stationkeeping mode (thruster attitude control) with all subsystems safe. GOES-10 continues to support the East satellite imaging and sounding operations. GOES-10 GVAR and all auxiliary services, including LRIT, DCS and EMWIN, are being broadcast simultaneously through GOES-10 and GOES-12. Those users monitoring GOES-12 for services may experience some dropouts during the engineering tests in the next few hours due to anticipated attitude disturbances. There could also be some signal interference at certain resulting from the simulcast transmission.


Red dot Due to the GOES-12 outage, the NOAA LRIT Broadcast and the NOAA DCS have been transferred to GOES-10 which is located at approximately 60 degrees west. This change will remain in effect until further notice or until the GOES 12 outage is resolved.


Red dot A GOES-12 North-South Station Keeping maneuver was performed on Tuesday, December 4, 2007 at 1756 UTC. An anomaly is currently ongoing. NOAA is not receiving GOES-12 (EAST) data until further notice. FULL DISK images are being captured on GOES-11 (WEST) until the problem is resolved. Data Affected by the Outage: GOES-12 (East) Imaging and Soundings data Date and Time of the Outage:12/4/2007, 1715 UTC, 12:15 PM, EST.


Red dot EUMETSAT will conduct a swap of Meteosat-9 and Meteosat-8: December 3-12, 2007. Products impacted will be all Meteosat-9 data and products from NOAA, including the Meteosat-9 DOMSAT Service. TheDate/Time of Initial Impact will be December 3, 2007 1200 UTC. Details can be found at:  http://www.eumetsat.int/Home/Main/Access_to_Data/index.htm?l=en  (link broken 11/16/2011).


Red dot Due to an operations oversight, the GOES-12 Reference Longitude was not changed as scheduled. It is now scheduled to be changed back to the nominal location of 75.0 degrees at J296/00:34UTC, 8:34PM EST tonight.


Red dot There was an LRIT outage on both operational GOES satellites on 9/22/2007 beginning from approximately 1400 UTC /10:00 AM EST for approximately 3 hours. The GOES imagery and GOES DCS parts of the LRIT service are again operational. The EMWIN stream on LRIT is still not up. We expect that it will be restored today, Monday, September 24th.


Red dot A set of EMWIN (Emergency Managers Weather Information) data is now available in the LRIT data stream. For more information on the EMWIN data please consult the EMWIN home page at http://www.weather.gov/emwin/index.htm and the LRIT section of the NOAASIS page at http://noaasis.noaa.gov/LRIT/.


Red dot This bulletin corrects the date of the Frequency Change Schedule to to 16 August. NOAA-18 VTX Frequency Change Schedule - The VTX on NOAA-18 will be switched to VTX-1 (137.10MHz). This will occur on Rev 11537 on 16 August 2007.


Red dot On 8 August, it was determined that NOAA-12 decommissioning is necessary since we are unable to regulate the bus power. Numerous attempts were made to preserve the mission, but the shunt degradation is not correctable. After more than 16 years of service, NOAA-12 will be decommissioned on Friday, 10 August, at 1710 UTC with the Fairbanks CDA contact. NOAA-12 was launched May 15, 1991.


Red dot Following an uncharacteristic anomaly on the Metop-A satellite, the Local Mission Direct Data Dissemination via HRPT will shortly be switched off and will remain off until further notice pending investigation. The anomaly occurred on day 2007.185 (July 4th) at approx. 1704 UTC, during orbit 03666.


Red dot The HIRS instrument noise level has quieted down and NOAA is ready for Satellite Operations Control Center (SOCC) to proceed with the test of APT transmitter effects on the HIRS instrument noise on NOAA 18. On 10 May, the following test scenario will be executed. Specific time period and orbits impacted by the test will be reported in a separate message. Proposed test scenario: The instrument will be placed in space view for two orbits in order to provide a uniform background and then we will systematically turn-on/off the various transmitters and monitor the noise looking for any noise spikes. After the command sequence has been run, the instrument will be returned to full-scan mode. APT user should expect to lose data from NOAA 18 during the testing.

On 10 May 2007, the HIRS test will be conducted on orbits 10154 and 10155.

Rev CDA Max El AOS FCT LCT LOS
10154 F 66.82 13:07:40 13:09:10 13:22:20 13:23:30
10155 F 27.92 14:48:20 14:50:10 15:02:30 15:03:00


Red dot On April 20 at 1014 UT, the MetOp spacecraft payload module shutdown all instruments. EUMETSAT is investigating and expects all instruments to remain off throughout the weekend. A review will be done on Monday, April 23. EUMETSAT expects to begin turning instruments back on by Monday afternoon (European time). A schedule will be provided. They expect to turn on AVHRR first and then AMSU-A.


Red dot "ESPC is not receiving any data on East and West LRIT receive systems for DCS data from GOES-11 GOES-12. The Command Data Acquisition site Wallops Island Virginia is having problems with the T-1 line and Domain 2. They are working on the problem. (4/19/07 0830 UTC) http://noaasis.noaa.gov/NOAASIS/ml/specialbull.html


Red dot "MetOp-A status is nominal except for the shut off of the LRPT service. Commissioning activities continue and are due to complete during May 2007 Due to commissioning activities service interruptions may occur between now and the middle of May. Satellite In-Orbit Validation (SIOV) completed successfully 2007-03-29." http://noaasis.noaa.gov/NOAASIS/ml/specialbull.html


Red dot Urgent Notice: After activation of the MetOp-A LRPT broadcast service in January a reduction in the quality of the HIRS instrument was recorded. After further testing it has been established by EUMETSAT that LRPT is the cause or contributing cause of the negative impacts. EUMETSAT has been investigating solutions, but with no viable option they announce the following to NOAA on February 5th 2007. "The initial investigations regarding the interference on MetOp-A HIRS have been completed, and the conclusion is that no scenario exists where LRPT on MetOp-A can be turned on without causing heavy interference on HIRS. Due to the operational importance of HIRS and the lack of an established LRPT user community, it is clear that LRPT will not be turned on again operationally on MetOp-A.". If users wish to submit comments or concerns please contact NOAA data services help desk at:  SPSD.Userservices@noaa.gov (address updated 11/16/2011).

Purpose of the LRPT broadcast service from the MetOp web site: "The digital LRPT service at VHF (137.1 MHz 72 kbps in QPSK) provides a subset of the HRPT data. It is comparable to the existing Automated Picture Transmission (APT) analogue service of the NOAA POES satellites. APT has, as a main objective, inexpensive access of low resolution Advanced Very High Resolution Radiometer (AVHRR) instrument images by local users; in America it is very popular in schools. MetOp augments the service by an innovative on-board compression scheme that allows LRPT transmission of AVHRR data at full instrument spatial and radiometric instead of providing low resolution images."


Red dot On January 24, 2007 EUMETSAT announced that the AHRPT Direct Readout service has been switched on again. The re-activation was successful with all telemetry parameters being nominal. The re-activation of AHRPT follows the successful LRPT Switch on performed on January 15, 2007. The AHRPT activation occurred on the 23 January 2007 at ca 08:30 UTC, using the nominal AHRPT frequency L-band 1701.3 MHz at 3.5 Mbps in QPSK. More information on MetOp-A AHRPT data can be found on the EUMETSAT web site:  http://www.eumetsat.int/Home/Main/Access_to_Data/Metop__amp__NOAA_Services/SP_1117714487826?l=en  (link broken 11/16/2011).


Red dot In the near future MetOp-A satellite will replace NOAA-17 as the primary operational morning satellite. In preparation for operational collection of MetOp data NOAA/NESDIS is holding an internal POES Data Constellation Meeting on January 25 to discuss changes to the current POES constellation. At the meeting the satellite staff will review the status for MetOp-A operations and the operational afternoon mission NOAA-18. In addition the meeting will examine what back-up satellite resources can be maintained with the addition of MetOp-A. NOAA's current back up satellites are NOAA-12, 14, 15, 16 , and eventually NOAA-17. Current thinking within NOAA is to decommission NOAA-14 which is currently in the same orbit plane as NOAA-17 and MetOp-A and creating issues with transmission interference with the ground. Direct Broadcast users of APT, and HRPT can send comments on their priorities to nesdis.osdpd.dsd.info@noaa.gov.


Red dot Notice: NOAA is investigating removing the Space Environment Monitor (SEM) instrument's data from the future High Data Rate (HRD) direct broadcast on the NPOESS Satellites. In addition SEM data would not be processed at NOAA's Satellite and Information Service, Environmental Satellite Processing Center (ESPC), but passed to a third party user, Space Weather Prediction Center (SWPC) of the National Weather Service (NWS) (Boulder, CO) for product generation. This would create a delay in access to SEM data for the direct broad cast community as products will need to be obtained from the third party rather then the spacecraft directly. If you have specific concerns please contact the Data Services Branch by e-mail at nesdis.osdpd.dsd.info@noaa.gov. Comments will be collect until November 28th 2006.


Red dot Wednesday October 18th NOAA 16 AVHRR scan motor current increased sharply over the weekend (October 14-15). NESDIS Satellite Services Division has reported a wavy pattern in the imagery ( as was before when this increased occurred). This situation is currently being monitored and any changes will be reported in a follow-up message.


Red dot On 10 Aug, (JD222) there will be a planned outage of data on NOAA-14 beginning around 14:30z, when there will be a spacecraft maneuver that will allow scientist to collect data that will characterize the side-lobes and any asymmetry of the Microwave Sounder Unit (MSU). Beginning with REV 59875 users can expect the outage to last several hours.


Red dot On 20 July 2006, on a contact beginning at 13:16z, NOAA engineers will change the APT frequency on NOAA-18 to use 137.9125MHz. In January 2006, in coordination with the direct readout group and EUM we temporarily switched to 137.1MHz due to the interference over Europe." Because of the pending Launch of Metop NOAA will be require to change frequency to avoid conflicts with the Metop spacecraft.


Red dot Fourth NOAA GOES-R Users’ Conference in Broomfield, Colorado, May 1-3, 2006. For more information see: http://www.ssd.noaa.gov/PS/SATS/MESS/MSG0461418.01.txt


Red dot NOAA-12 and NOAA-15 will enter a conflict period requiring NOAA-15 APT to be turned OFF from March 1, 2006 at 1800 UTC until April 2, 2006 at 2300 UTC. NOAA-12 APT will continue to transmit during the conflict period.


Red dot Update: NOAA-18 APT frequency change on January 04, 2006 at 0903 UTC - NOAA implemented a trial frequency change on NOAA-18 APT transmissions from 137.9125 MHz (VTX-2) to 137.10 MHz (VTX-1) on January 4, 2006 while the spacecraft was over the South Pole on revolution #3227. NOAA will re-evaluate this frequency change following the MetOp-A launch scheduled for June 30, 2006 as part of the U.S.A. and EUMETSAT cooperative Initial Joint Polar System (IJPS) agreement.


Red dot NOAA-N successfully launched on May 20, 2005 and was declared operational on August 30, 2005 as NOAA-18.


Red dot NOAA polar satellite TBUS bulletins updated to re-introduce NOAA-12 and 14 on August 29. Data will be valid for September 1, 2005.

TBUS1 NOAA-17 sat id 37
TBUS2 NOAA-14 sat id 38 - back to previous position
TBUS3 NOAA-15 sat id 34
TBUS4 NOAA-16 sat id 36
TBUS5 GOES-10
TBUS6 GOES-12
TBUS7 NOAA-18 sat id 39
TBUS8 NOAA-12 sat id 43


Red dot Update: Visit the 2004 Satellite Direct Readout Conference web site for conference proceedings and post conference information -  The Conference was held on December 6-10, 2004 in Miami, Florida and it was a huge success with over 200 participants from 29 countries in attendance. http://directreadout.noaa.gov/miami04/


Red dot LRIT is now transmitted full time on GOES East and GOES West. Get the latest news on LRIT from the > LRIT Homepage


Red dot GOES East and GOES West Eclipse > Schedules


Red dot GOES Single Chord Operations > SINGLECH

Also see the GOES Special Bulletins page at the Satellite Services Division web site


SOCC morning report updated at approximately 1500 UTC, Monday through Friday (secondary location).

NOAA/NESDIS/OSDPD; Page Reviewed/Revised Daily

 

Satellite Products and Services Division
Direct Services Branch
Phone: 301-817-4521/4523
Take Our Survey
Mailing Address:
NSOF (E/SPO53)
1315 East-West Hwy
Silver Spring, MD 20910-3282     USA